In a triangle, the middle side is 7cm longer than the shortest side. The longest side is 8cm more than the shortest side. The perimeter is 31cm. What is the length of each side of the triangle? Must show work! thanks
Answer by scott8148(6628)

let x=(shortest side), so x+7=(middle side) and x+8=(longest side)

perimeter=x+(x+7)+(x+8) ... 31=3x+15 ... 16=3x ... x=5 1/3

x+7=12 1/3 ... x+8=13 1/3